OpenStudy (anonymous):

How many molecules of carbon dioxide are there in 44.0 grams of the gas? (atomic mass of carbon = 12.0 and atomic mass of oxygen = 16.0)

OpenStudy (unklerhaukus):

The molecular mass \(M\) is the sum of atomic masses of the atoms in the molecule. For carbon dioxide; CO_2, \[M_{\text{CO}_2}=M_\text C+2M_\text O\\\qquad\quad=12.0\,[\tfrac{\text g}{\text{mol}}]+2\times16.0\,[\tfrac{\text g}{\text{mol}}]\\%\qquad\quad=44\,[\tfrac{\text g}{\text{mol}}]\] Use the molecular mass of CO_2 to convert mass, to a number of moles \(n\) \[M\triangleq \frac mn\implies n=\frac mM\] Avogadro's number \(N_A\) is the number of particles \(N\) in one mole of any substance, so simply multiply the number of moles \(n\) you found above by Avogadro's number \[N_A=6.022\times10^{23}\,[\text{particles}/\text{mol}]\] \[N=nN_A\]